OpenStudy (anonymous):

Select the pair of equations whose graphs are perpendicular. A. 2y = –3x + 5 2x + 3y = 4 B. y = 2x – 7 x + 2y = 3 C. 5x – 8y = 9 12x – 5y = 7 D. x + 6y = 8 y = 2x – 8

OpenStudy (anonymous):

@jim_thompson5910 @BlueEyedCountryGirl

jimthompson5910 (jim_thompson5910):

hint: two lines are perpendicular if their slopes multiply to -1 example: y = 2x + 3 y = -1/2x + 7 are perpendicular because 2*(-1/2) = -1

OpenStudy (anonymous):

@jim_thompson5910 okay so what's the answer? Not trying to be rude.. haha

jimthompson5910 (jim_thompson5910):

sorry I can't just give out the answers, but I can help you get there

jimthompson5910 (jim_thompson5910):

are you able to solve for y?

OpenStudy (anonymous):

Yeah well can you work through it with me?

jimthompson5910 (jim_thompson5910):

how would you solve 2y = –3x + 5 for y?

jimthompson5910 (jim_thompson5910):

how would you move that 2 over?

OpenStudy (anonymous):

that's the part I don't get...

OpenStudy (anonymous):

I mean I don't get the whole thing but this is also a problem.

jimthompson5910 (jim_thompson5910):

well 2y means 2 times y to move it over, you have to undo multiplication...so you have to divide both sides by 2

jimthompson5910 (jim_thompson5910):

\[\Large 2y = –3x + 5\] \[\Large y = \frac{–3x + 5}{2}\] \[\Large y = \frac{–3x}{2} + \frac{5}{2}\] \[\Large y = -\frac{3}{2}x + \frac{5}{2}\] What is the slope here?

OpenStudy (anonymous):

I honestly have no idea... sorry. I have a major gap from 5th grade till now.

jimthompson5910 (jim_thompson5910):

you're familiar with y = mx+b right? or no?

OpenStudy (anonymous):

sorry... no *winces*

jimthompson5910 (jim_thompson5910):

that's ok, y = mx+b is the general form of any line where m is the slope and b is the y-intercept

jimthompson5910 (jim_thompson5910):

the slope determines how steep/shallow the line is the y-intercept moves the line up and down so these two uniquely determine a line

jimthompson5910 (jim_thompson5910):

compare \(\Large y = -\frac{3}{2}x + \frac{5}{2}\) to y = mx+b what is m in this case?

OpenStudy (anonymous):

-3/2 ??

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now let's solve 2x + 3y = 4 for y

OpenStudy (anonymous):

Yayy! I'm honestly so proud right now. :)

jimthompson5910 (jim_thompson5910):

That's good. I'm glad you are. Let's solve the second equation for y. \[\Large 2x + 3y = 4\] \[\Large 3y = 4 - 2x\] \[\Large 3y = -2x + 4\] \[\Large y = \frac{-2x + 4}{3}\] \[\Large y = \frac{-2x}{3} + \frac{4}{3}\] \[\Large y = -\frac{2}{3}x + \frac{4}{3}\] What is the slope here?

OpenStudy (anonymous):

I...don't know :/

jimthompson5910 (jim_thompson5910):

what is m in this case?

jimthompson5910 (jim_thompson5910):

go back to y = mx+b

OpenStudy (anonymous):

okay give me a min.

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

is it 2???

jimthompson5910 (jim_thompson5910):

compare \(\Large y = -\frac{2}{3}x + \frac{4}{3}\) to y = mx+b

OpenStudy (anonymous):

oh so -2/3 ?

jimthompson5910 (jim_thompson5910):

yep, so the two slopes of the two equations are -3/2 and -2/3

jimthompson5910 (jim_thompson5910):

what do they multiply to?

OpenStudy (anonymous):

so multiply those?

OpenStudy (anonymous):

-1? or 1?

jimthompson5910 (jim_thompson5910):

you have two negative numbers multiplied

jimthompson5910 (jim_thompson5910):

so is the result negative? or positive?

OpenStudy (anonymous):

1. Positive 1.

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

since the result is NOT -1, that means the two lines are NOT perpendicular

OpenStudy (anonymous):

Oh okay..

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

I have ten mins to finish this test....